OpenStudy (anonymous):

write the quadratic equation whose roots are 5 and 6 and whose leading cofficient is 4

OpenStudy (anonymous):

can someone help me please

OpenStudy (anonymous):

Ok. let the equation be ax^2 + bx+c. you told leading coefficient a = 4, See that sum of roots = -b/a, product of roots = c/a. Now try out. :)

OpenStudy (anonymous):

@luckythebest WAHT IS A AND B

OpenStudy (anonymous):

a is the leading coefficient i.e. in ax2, and B is the coefficient of x i.e. bx ax2+bx+c, where sum of roots = -b/a and product of roots = c/a. a = 4. ( as given)